OpenStudy (anonymous):

Factor the expression: 2x2 + 5x + 3

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

@chosenmatt

OpenStudy (chosenmatt):

which number you think it is -2 or -1

OpenStudy (anonymous):

-1

OpenStudy (chosenmatt):

yup good job

OpenStudy (anonymous):

what do i do now then @chosenmatt

OpenStudy (chosenmatt):

x=-1 what you just said so plug it in

OpenStudy (anonymous):

2*-1*2+5*-1+3

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

@mathmath333

OpenStudy (mathmath333):

first u have to find the factors of \(\Large ax^2+bx+c\) \(\Large a \) and \(\Large c \) here \(\Large 2x^2+5x+3\\ \) \(\Large 2\times 3=6\) therefore \(\Large 6\) has factors \(\Large 2\) and \(\Large 3\) also \(\Large 2\) and \(\Large 3\) add to \(\Large 5\) which is the \(\Large b\) term hence u can write \(\Large b\) term which is \(\Large 5\) as \(\Large 2+3\) therefore \(\Large 2x^2+5x+3\\ =\Large 2x^2+2x+3x+3\\ =\Large 2x(x+1)+3(x+1)\\ =\huge (x+1)(2x+3)\\\)